# Ultra-Fast Language Generation via Discrete Diffusion Divergence Instruct

### v2

@@ 5.2 Ablation Studies
We conduct comprehensive ablation studies to validate the contribution of each component in DiDi-Instruct. We perform two types of analyses: a cumulative study (Table 2) that progressively adds techniques to a baseline, showing their synergistic benefits, and a leave-one-out study (Table 2) that removes individual components to confirm their necessity. Implementation details can refer to F.3.
@@
Our ablation studies reveal distinct roles for each component in our framework. We find that a two-step score decomposition is a non-negotiable cornerstone, providing essential stability without which the model fails to train. The most significant performance gains are driven by coupling $t$ in (5) and (9). Loss shaping with $\omega(t)$} and $\pi(t)$} further smooths optimization (especially around 16 NFEs, while effects at very small/large budgets are modest). Finally, we identify two budget-dependent components: Regularization is crucial for stability at very few NFEs ($\leq 8$ NFEs) but detrimental at higher budgets, while Guided Inference boosts quality at low NFEs and enhances diversity at high NFEs. These findings highlight a hierarchy of importance and nuanced interactions between the techniques. More details are presented in Appendix F.4.
5.3 Model Scaling Up
We perform an incremental scaling experiment to 424M parameters while holding the training and inference pipeline fixed to the 169M configuration. The teacher model is a pre-trained 424M MDLM, which is a Diffusion Transformer with a hidden size of 1024, 24 blocks, and 16 attention heads (check Table 4 for detailed network structure). The student is distilled on a single H100 using the same data, masking policy, optimizer, and schedule. To balance the training, we also scale the discriminator to 373M and employ a deeper network with more trainable parameters.
The generative PPL and entropy are shown in Figure 4, and we report improvements relative to the teacher at matched NFEs here: from 8 to 1024 NFEs, DiDi-Instruct yields large and consistent PPL reductions of $88.5\%$ (8 NFEs), $87.7\%$ (16 NFEs), $85.3\%$ (32 NFEs), $82.7\%$ (64 NFEs), and $79.0\%$ (128 NFEs). Entropy remains comparable throughout, indicating that diversity is preserved while accuracy improves. These results confirm that the quality–efficiency advantages of DiDi-Instruct persist at increasing scale with minimal procedural changes. For more details, please refer to Appendix F.5.
@@ 5.4 Another Application: Protein Sequences Generation
Our results demonstrate that the distilled student consistently achieves superior pLDDT scores across generation settings ranging from 8 to 512 NFEs. Compared to the teacher model, the student not only preserves the ability to generate variable-length protein sequences but also enhances structural quality in most cases. Moreover, our model surpasses the high-confidence threshold ($\text{pLDDT}>70$) with as few as 8–32 NFEs, while the teacher requires substantially more NFEs to reach a comparable level. These results highlight that our distillation framework not only improves structural confidence but also delivers stable performance across sequence lengths and generation budgets. For experimental details on protein sequence distillation, please refer to Appendix F.6, where we also provide visualizations of low-confidence protein sequences generated by DPLM under fewer NFEs in Figure 6, along with high-quality examples generated by DiDi-Instruct across different NFEs in Figure 7.
@@ 6 Conclusion and Future Work
In this work, we introduced DiDi-Instruct, a training-based acceleration framework for fast language generation that distills a high-quality teacher into a few-step student. Our design targets three axes simultaneously: (i) objective design via a tractable policy–gradient update driven by a discriminator–estimated reward; (ii) training stability through score decomposition and grouped reward normalization; and (iii) inference efficiency through reward-guided ancestral sampling with gradient tilting and re-ranking. Experiments demonstrate strong gains in generation quality, large reductions in training/inference time, and competitive zero-shot generalization, corroborated by comprehensive cumulative, leave-one-out ablations, model scaling up, and protein sequence generation.
@@
We plan to scale DiDi-Instruct to billion-parameter models, which presents a practical challenge due to the GPU memory requirements of concurrently maintaining the teacher, student, and discriminator. Nevertheless, our findings already establish a new state-of-the-art trade-off among comparable methods, with the student model excelling in quality at low NFEs and maintaining diversity at higher computational budgets. We posit that DiDi-Instruct offers a foundational recipe (principled objectives, training stability, and efficient guidance) for developing high-performance generative models.
Appendix A Language Generation Showcase
To provide a qualitative illustration of our method’s effectiveness, we present a series of uncurated samples from the base and distilled model below. The first sample is generated from the 1024-step teacher model, serving as a high-quality reference. Subsequent samples are generated by DiDi-Instruct, with an increasing NFEs from 8 to 128. These examples clearly demonstrate the progressive improvement in coherence and semantic quality as the sampling budget increases, showcasing how DiDi-Instruct efficiently approaches the teacher’s performance. For more examples, please refer to Appendix F.7.
Samples from model obtained with $1024$ NFEs teacher model. Preplexity=$38.53$, Entropy=$5.22$. He parked and it was then that I saw that president and ended up there, speaking with a journalist. The Prince was granted anonymity and noted that the FBI was aware of the familiarity with Russia and was conducting the investigation as appropriate. Victor Atkinson, an FBI chief-officer who is retired, hacked the personal PINs and cell phone lines of the two Trump officials to conceal the information, as shown inside the countrys electronic networks. Anonymous messages can appear in our apps like Twitter and Facebook.
Samples from model distilled by DiDi-Instruct with 88 NFEs. Preplexity=$62.24$, Entropy=$5.17$. In what some call a landmark technical proposal toward launching a space shuttle, Intel is set to make the flight of Americas first space shuttle happen next week. Starting next week aboard the Space Launch System, the waveS memory mission will be a mission to use Intels wave memory technology. This technology is called vertical wave memory, whose idea was first conceived earlier this week. "While it remains not yet seen whether the mission will take place on Americas first space flight, it is a noting that the idea was first conceived earlier this week," President Barack Obama said.
Samples from model distilled by DiDi-Instruct with $16$ NFEs. Preplexity=$30.99$, Entropy=$5.22$. The U.S. the US Senate on Tuesday introduced a bill that would allow crocodiles of adult sons and their daughters to paddle lasers into the range of eyes and ears. However, it will require lawmakers from the Senate House of Representatives to ask questions in a bill version of blinding optics. An aide from the House Majority Whip, the Senate Republican Whip, introduced a bill that would require a hard drive in the eyes of adult sons and their daughters. It will require access from lawmakers to a hard drive to ask questions in a bill known as blinding optics.
Samples from model distilled by DiDi-Instruct with $32$ NFEs. Preplexity=$23.60$, Entropy=$5.20$. Intel today is taking back the high-end of the PC market, and its sooner rather than later, magazine Pocket reports. The high-end processor is taking advantage of the high specs, and the renowned chip firm is no exception. The Intel processor is sporting a 2nd wave memory clocked. This is up to a total of 3GB of memory bandwidth and 4GB of GDDR5. The 2nd memory is available for 999 pounds and is available in June. The memory is also 18 times faster than a PC with the same amount of GDDR5.
Samples from model distilled by DiDi-Instruct with $64$ NFEs. Preplexity=$19.61$, Entropy=$5.18$. PENNAI, China - IBM and Hewlett-Pack expressed hopes over docking in space companies next-generation computers during a joint press conference on Dec. 5. The companys chairman Chongq Iardai has developed a docking system to improve space companies next-generation computer systems. This was according to the exclusive interview published by Forbes magazine. A consortium of Hewlett-Packard, including Intel and IBM has created a new wave of docking systems, which could improve the companys current-generation computer systems.
Samples from model distilled by DiDi-Instruct with $128$ NFEs. Preplexity=$17.50$, Entropy=$5.17$. In its first foray into market that created top-end graphics, Intel announced it plans to launch a next-generation graphics chip that is responsible for powering the personal computers power system. The announcement was made Monday during a joint press conference held by Las Vegas-based Intel. The conference also included the Florida-based technology company Ion Graphics. The new graphics chip, which is slated for a 2017 release, powers the Windows operating system. Ion Graphics is the worlds largest graphics company, launching its initial public offering in the 2000’s.
Appendix B Extended Related Works
B.1 Related Work on Continuous-Time Diffusion Models
@@
Diffusion models (Song et al., 2021b; Ho et al., 2020; Sohl-Dickstein et al., 2015) added Gaussian noises to data with a forward diffusion process and then learn the backward stochastic differential equation (SDE) with so-called score neural networks. Early acceleration reduced the steps of iterative samplers using training-free high-order ODE solvers (Song et al., 2021a; Lu et al., 2022; Karras et al., 2022; Xue et al., 2023). However, solver-based acceleration suffered from discretization error in less than 10 generation steps.
@@
To achieve extremely few-step generative modeling, researchers have studied various ways of diffusion distillation (Luo, 2023). Luhman and Luhman (2021) and Salimans and Ho (2022) first introduce training the few-step student models by learning the consecutive trajectory mapping of diffusion solvers. The seminal work of Song et al. (2023) and Luo et al. (2023b) opens the one-step generation of diffusion models through different distillation principles. Specifically, consistency models (Song et al., 2023) distill the few-step models with the trajectory consistency principle, resulting in strong performances (Song and Dhariwal, 2024; Lu and Song, 2025; Geng et al., 2025b; Luo et al., 2023a; Kim et al., 2024; Geng et al., 2025a). Diff-Instruct (Luo et al., 2023b) introduces the distribution matching principle that distills one-step generative models, resulting in leading efficient image generative models (Wang et al., 2025; Luo et al., 2025a, 2024; Luo, 2025; Xu et al., 2025; Yin et al., 2024b, a; Xie et al., 2024; Fan et al., 2023; Zhou et al., 2024; Huang et al., 2024; Luo et al., 2025b, c).
Many other works have also studied the few-step continuous-space generative models from different perspectives (Liu et al., 2022; Geng et al., 2023; Zhou et al., 2025; Nguyen and Tran, 2023; Lin et al., 2024; Boffi et al., 2024; Xu et al., 2023; Xiao et al., 2021; Meng et al., 2022; Zhang and Chen, 2022; Sauer et al., 2023; Ren et al., 2024; Yan et al., 2024; Gu et al., 2023; Chen et al., 2024).
B.2 Related Work on Discrete-Time Diffusion Models
@@
Early attempts to model discrete data with diffusion-framed categorical transitions via multinomial/argmax flows (Hoogeboom et al., 2021). Later, it generalized them with Structured Denoising Diffusion Models (D3PM) and introduced structured transition matrices (e.g., discretized Gaussian kernels, nearest-neighbor matrices, and absorbing states) and an auxiliary cross-entropy loss (Austin et al., 2021). Continuous-time viewpoints and relaxations further clarified the link between discrete corruption processes and stochastic dynamics (Campbell et al., 2022; Dieleman et al., 2022; Chen et al., 2023). For language generation, diffusion has been explored for controllable text and sequence-to-sequence tasks (Li et al., 2022b; Gong et al., 2023).
A line of work centers on MDMs for language modeling: Lou et al. (2024) cast discrete diffusion training as estimating time-indexed ratios of the data distribution, which provides a complementary alternative to score- and cross-entropy–based objectives. Sahoo et al. (2024) showed that simple training recipes (including a Rao-Blackwellized objective) can approach auto-regressive perplexity, while Shi et al. (2024) derived a weighted time-integral of cross-entropy losses that enables state-dependent masking schedules and improves downstream quality. Recent theory deepens our understanding of absorbing processes: the absorbing score admits a conditional-probability interpretation and yields time-independent formulations that connect diffusion with any-order auto-regressive models (Ou et al., 2025); in parallel, Zheng et al. (2025) proved an equivalence between masked diffusion and time-agnostic masked modeling and highlighted numerical issues in categorical sampling.
Scaling and inference efficiency are active topics. Large-parameter diffusion language models achieve competitive performance with strong auto-regressive LLMs (Nie et al., 2025), and adaptive token ordering shows that planning which positions to denoise can markedly improve reasoning (Kim et al., 2025). On the sampling side, informed correctors learn predictor–corrector schemes (Zhao et al., 2024), schedule optimization mitigates compounding decoding error (Park et al., 2024), and plan-and-denoise approaches add a planner that selects the most corrupted positions to improve both efficiency and quality (Liu et al., 2025).
Building on these advancements, our work introduces a discrete diffusion distillation framework based on distribution matching. We aim to enhance the performance of discrete diffusion models for few-step generation, thereby enabling faster and higher-quality text generation.
B.3 Related Work on Few-Step Discrete Diffusion LLMs
@@
To distill dLLMs into few-step generators for fast language generation, a natural approach is to extend the well-developed distillation techniques already established for continuous diffusion models to the discrete domain. However, in MDMs, all tokens are eventually mapped to the masked state, meaning the prior distribution collapses to the fully masked sequence. As a result, generation under MDMs is inherently deterministic, in contrast to the stochastic reverse dynamics in continuous diffusion. Consequently, the stochasticity in MDMs arises from the choice of the masking schedule; once a schedule is fixed, the reverse path is deterministic. This stands in sharp contrast to the inherently stochastic reverse dynamics of continuous diffusion. This determinism is problematic for distillation, as training a student model solely on these fixed trajectories risks mode collapse and a failure to capture the teacher’s generalization capabilities. This fundamental difference makes the extensive body of work on continuous-state diffusion distillation not directly applicable. To make this distinction concrete, we first show how the following methods address these challenges in discrete diffusion distillation, followed by a comparison of these representative approaches.
@@
SDTT (Deschenaux and Gulcehre, 2025). SDTT progressively distills a multistep teacher diffusion into a fewer-step student model. Due to the absence of an ODE trajectory as in continuous cases, it matches the teacher-student distribution via KL divergence minimization. Similar to DUO (Sahoo et al., 2025), it needs to decrease the number of student steps for training stability gradually. Such a curriculum-based strategy can be inefficient and computationally costly.
@@
DUO (Sahoo et al., 2025). Continuous-space consistency distillation (CD) requires probability flow ODE (PF-ODE) trajectories, while MDMs are inherently stochastic, making direct CD inapplicable. DUO overcomes this by connecting Gaussian diffusion with USDMs, enabling CD in the discrete setting. By applying an $\operatorname*{arg\,max}$ to the latent vectors of Gaussian diffusion, continuous vectors are mapped into discrete one-hot tokens. It is proven that the resulting marginal distribution exactly matches that of a USDM under a suitably transformed noise schedule. Consequently, the evolution of a USDM can be described by an ODE, which makes CD feasible. However, performing CD on USDM requires multiple rounds of distillation and a carefully annealed discretization schedule for training stability. Specifically, the CD procedure starts from the USDM diffusion and gradually learns to map two points, $t$ and $t-\Delta t$, along the same ODE trajectory back to the same origin. Initially, since the diffusion model cannot take large steps without losing accuracy, $\Delta t$ must be kept small. Once the model learns to consistently align $t$ and $t-\Delta t$, the step size can be increased, eventually reaching the largest possible step $\Delta t=t$, i.e., the distilled model can jump large steps for few-step inference. A small step is safe but biased, while a large step is unbiased but unstable. This tradeoff is similar to the observation of CD in the continuous domain.
@@
DiMO (Zhu et al., 2025b). DiMO distills multi-step MDMs into a one-step generator. The key idea is to augment the prior distribution: instead of restricting the initial state to a fully-masked sequence, DiMO samples a subset of tokens from the entire vocabulary. This relaxation enriches the prior, allowing the model to leverage partial information during generation and improving both efficiency and diversity of outputs. During training, DiMO employs an on-policy distillation strategy. The one-step generator is supervised to match the teacher’s conditional prediction distribution, not its final output distribution. Specifically, it uses the generator’s own one-step output to create a pseudo-intermediate state, and then minimizes the divergence between the student’s and the teacher’s predicted token distributions conditioned on this state. To avoid mode collapse and reduce mismatch with the teacher’s training distribution, DiMO introduces a token initialization strategy that mixes mask tokens with random tokens and adds Gaussian perturbations to embeddings.
@@
Advantages of DiDi-Instruct. Our work presents a straightforward and intuitive framework for distilling dLLMs by directly matching teacher and student distributions, addressing several key limitations of recent approaches. Unlike multi-stage progressive distillation techniques (Sahoo et al., 2025; Deschenaux and Gulcehre, 2025), our method is a direct, single-stage process. It not only avoids the use of deterministic or heuristic mappings, but also dispenses with the compute-intensive, multi-round training used by consistency-based approaches, which requires careful choice of time discretization step annealing. Furthermore, by aligning the student with the teacher’s entire stochastic process, our framework offers broad applicability to general dLLMs, in contrast to methods tailored for specific architectures, e.g., DUO (Sahoo et al., 2025) relies on USDM for duality. Finally, we provide a mathematically rigorous solution to the problem of non-differentiability in discrete spaces. This avoids the need for biased proxy-gradient estimators, whose performance in purely textual domains remains underexplored (Zhu et al., 2025b), and instead offers a principled path for distillation.
Appendix C Student Objective
Our aim is to train a masked-diffusion student whose forward-time marginals match those of a teacher across the entire time horizon. We adopt the IKL between the student and teacher forward marginals $\mathbf{q}_{\nu}$ and $\mathbf{q}_{\theta}$ as the training objective. In the masked setting, the forward-time corruption kernel is a forward absorbing process, e.g.,

### v3

@@ 4.2 Ablation Studies
We conduct comprehensive ablation studies to validate the contribution of each component in DiDi-Instruct. We perform two types of analyses: a cumulative study (Table 2) that progressively adds techniques to a baseline, showing their synergistic benefits, and a leave-one-out study (Table 2) that removes individual components to confirm their necessity. Implementation details can refer to E.5.
@@
Our ablation studies reveal distinct roles for each component in our framework. We find that a two-step score decomposition is a non-negotiable cornerstone, providing essential stability without which the model fails to train. The most significant performance gains are driven by coupling $t$ in (5) and (9). Loss shaping with $\omega(t)$} and $\pi(t)$} further smooths optimization (especially around 16 NFEs, while effects at very small/large budgets are modest). Finally, we identify two budget-dependent components: Regularization is crucial for stability at very few NFEs ($\leq 8$ NFEs) but detrimental at higher budgets, while Guided Inference boosts quality at low NFEs and enhances diversity at high NFEs. These highlight a hierarchy of importance and nuanced interactions between the techniques.
To further explore RGAS, we analyze the hyperparameters $h$ and $M$ through performance landscape scans and Functional ANOVA (Hutter et al., 2014). Our analysis reveals that at low NFEs (<16<16), $h$ dominates performance by guiding global structure; as the computational budget increases (e.g., 32 NFEs), the importance of $M$ rises significantly. More details are presented in Appendices E.5-E.7.
4.3 Downstream Task Evaluation
To validate the practical utility, we evaluate DiDi-Instruct and baselines on downstream tasks. Specifically, we conduct experiments on Domain Adaptation (where we fine-tune models on the MMLU (Hendrycks et al., 2021) and PubMed benchmarks), and Frozen Feature Extraction (using the GLUE MRPC dataset (Dolan and Brockett, 2005)).
On MMLU fine-tuning, DiDi-Instruct achieves the largest reduction in negative log-likelihood after 5,000 fine-tuning steps while matching teacher accuracy. On PubMed, DiDi-Instruct nearly matches the teacher’s perplexity while outperforming other distillation methods. Moreover, in frozen-feature evaluation on GLUE MRPC, DiDi-Instruct achieves the best accuracy and F1 score, demonstrating superior semantic representation quality. These results confirm that DiDi-Instruct maintains strong downstream performance while providing substantial efficiency gains. See Appendix E.4 for details.
4.4 Scaling Up Teacher Model Size
We conduct an incremental scaling study to 424M parameters while keeping the training and inference pipeline identical to the 169M configuration. The teacher model is a pre-trained 424M MDLM, which is a DiT-style transformer with hidden size 1024, 24 blocks, and 16 attention heads (See Table 3 for network architecture details). The student is distilled on a single H100 using the same data, masking policy, optimizer, and schedule. To balance the training, we also scale the reward discriminator to 373M parameters by adopting a deeper network with more trainable parameters.
Figure 4 reports generative PPL and entropy. We report improvements relative to the teacher at matched NFEs. From 8 to 1024 NFEs, DiDi-Instruct yields large and consistent PPL reductions relative to the teacher: $88.5\%$ (8 NFEs), $87.7\%$ (16 NFEs), $85.3\%$ (32 NFEs), $82.7\%$ (64 NFEs), and $79.0\%$ (128 NFEs). Notably, with only 16 NFEs, the distilled model reaches a perplexity of 32.79, marking an 11.4% improvement over the 1024-step teacher baseline. Entropy remains comparable under the same NFE budgets, indicating that diversity is preserved while accuracy improves. Overall, these results confirm that the quality–efficiency advantages of DiDi-Instruct persist at a larger scale with minimal procedural changes. Additional details are provided in Appendix E.8.
@@ 4.5 Another Application: Protein Sequences Generation
Our results demonstrate that the distilled student consistently achieves superior pLDDT scores across generation settings ranging from 8 to 512 NFEs. Compared to the teacher model, the student not only preserves the ability to generate variable-length protein sequences but also enhances structural quality in most cases. Moreover, our model surpasses the high-confidence threshold ($\text{pLDDT}>70$) with as few as 8–32 NFEs, while the teacher requires substantially more NFEs to reach a comparable level. These results highlight that our distillation framework not only improves structural confidence but also delivers stable performance across sequence lengths and generation budgets. Appendix E.9 includes detailed setup, and visual comparisons between low-confidence outputs of DPLM at small NFEs and high-quality samples produced by DiDi-Instruct (Figures 6-7).
To ensure that the observed improvements in structural confidence (pLDDT) do not stem from mode collapse, we evaluate sequence diversity using MMseqs2 clustering (Steinegger and Söding, 2017). Quantitative results confirm that DiDi-Instruct maintains competitive cluster entropy and low cluster sizes comparable to the teacher model, particularly in few-step regimes. This indicates that our method generates diverse, biologically meaningful sequences without collapsing to a few high-confidence modes. Detailed results are provided in Appendix 7.
@@ 5 Conclusion and Future Work
In this work, we introduced DiDi-Instruct, a training-based acceleration framework for fast language generation that distills a high-quality teacher into a few-step student. Our design targets three axes simultaneously: (i) objective design via a tractable policy–gradient update driven by a discriminator–estimated reward; (ii) training stability through score decomposition and grouped reward normalization; and (iii) inference efficiency through reward-guided ancestral sampling with gradient tilting and re-ranking. Experiments demonstrate strong gains in generation quality, large reductions in training/inference time, and competitive zero-shot generalization, corroborated by comprehensive cumulative and leave-one-out ablations.
@@
We plan to scale DiDi-Instruct to billion-parameter models, which presents a practical challenge due to the memory requirements of concurrently maintaining the teacher, student, and discriminator. Nevertheless, our findings already establish a new state-of-the-art trade-off among comparable methods, with the student model excelling in quality at low NFEs and maintaining diversity at higher computational budgets. We posit that DiDi-Instruct offers a foundational recipe (principled objectives, training stability, and efficient guidance) for developing high-performance generative models.
Acknowledgments
We thank the CoreWeave AI cloud platform for supporting part of the computational resources used in this work. Nan Jiang acknowledges support from the Texas Advanced Computing Center (TACC) under award CCR25054. Guang Lin acknowledges support from the National Science Foundation (NSF) under grants DMS-2533878, DMS-2053746, DMS-2134209, ECCS-2328241, CBET-2347401, and OAC-2311848; the U.S. Department of Energy (DOE) Office of Science, Advanced Scientific Computing Research program under award DE-SC0023161; the SciDAC LEADS Institute; and the DOE Fusion Energy Sciences program under grant DE-SC0024583.
Appendix A Related Works
A.1 Related Work on Continuous-space Diffusion Distillation
@@
Diffusion models (Song et al., 2021b; Ho et al., 2020; Sohl-Dickstein et al., 2015) perturb data by adding Gaussian noise in a forward diffusion process and then learn the reverse-time dynamics (formulated as a stochastic differential equation) using score networks. Early efforts to accelerate sampling reduced the number of calls to iterative samplers using training-free high-order ODE solvers (Song et al., 2021a; Lu et al., 2022; Karras et al., 2022; Xue et al., 2023). However, such solver-based acceleration still suffered from discretization error in less than ten generation steps.
@@
To achieve few-step generation, Luhman and Luhman (2021) and Salimans and Ho (2022) first introduce training the few-step student models by learning the consecutive trajectory mapping of diffusion solvers. Subsequently, the seminal work of Song et al. (2023) and Luo et al. (2023b) opens the one-step generation of diffusion models through different distillation principles. Specifically, consistency models (Song et al., 2023) distill the few-step models with the trajectory consistency principle, resulting in strong performances (Song and Dhariwal, 2024; Lu and Song, 2025; Geng et al., 2025b; Luo et al., 2023a; Kim et al., 2024; Geng et al., 2025a). Diff-Instruct (Luo et al., 2023b) introduces the distribution matching principle that distills one-step generative models, resulting in leading efficient image generative models (Wang et al., 2025; Luo et al., 2025a; 2024; Luo, 2025; Xu et al., 2025; Yin et al., 2024b; a; Xie et al., 2024; Fan et al., 2023; Zhou et al., 2024; Huang et al., 2024; Luo et al., 2025b; c). Later, many other works have also studied the few-step continuous-space generative models from different perspectives (Liu et al., 2023; Geng et al., 2023; Zhou et al., 2025; Nguyen and Tran, 2024; Lin et al., 2024; Boffi et al., 2025; Xu et al., 2024; Xiao et al., 2021; Meng et al., 2023; Zhang and Chen, 2023; Sauer et al., 2023; Ren et al., 2024; Yan et al., 2024; Gu et al., 2024; Chen et al., 2024).
Inspired by Luo et al. (2023b), this work leverages the distribution-matching principle to tackle the more challenging problem of discrete language sequence generation.
A.2 Related Work on Discrete Diffusion Models
@@
Early studies of discrete diffusion modeled categorical data via multinomial or argmax-based transitions (Hoogeboom et al., 2021). This line of work was later generalized by D3PM, which introduced structured transition matrices (such as discretized Gaussian kernels, nearest-neighbor transitions, and absorbing states) together with an auxiliary cross-entropy objective (Austin et al., 2021). Continuous-time perspectives and relaxations further clarified the connection between discrete corruption processes and stochastic dynamics (Campbell et al., 2022; Dieleman et al., 2022; Chen et al., 2023). In parallel, diffusion-based ideas were also explored for controllable text generation and sequence-to-sequence modeling (Li et al., 2022; Gong et al., 2023).
For language modeling, two principal families of dLLMs have emerged. MDMs (Lou et al., 2024; Sahoo et al., 2024; Shi et al., 2024; Ou et al., 2025) treat the forward process as progressive masking of tokens. At the highest noise level, every token is replaced by a special [MASK] symbol; the reverse process gradually unmask tokens. Once a token is unmasked, it remains fixed, but many tokens can be denoised simultaneously. Intuitively, MDMs essentially train BERT (Devlin et al., 2019) under a hierarchy of noise levels, motivated by a scaling-based rationale for generation (He et al., 2023; Sahoo et al., 2024).Uniform‑state diffusion models (USDMs) (Austin et al., 2021; Gulrajani and Hashimoto, 2023; Schiff et al., 2025; Sahoo et al., 2025) instead corrupt tokens by sampling from a uniform distribution over the vocabulary. Consequently, each token can change multiple times during sampling, enabling self‑correction and strong theoretical connections to continuous Gaussian diffusion.
Additionally, recent work on dLLMs has advanced along several fronts, including training objectives and theoretical foundations (Lou et al., 2024; Sahoo et al., 2024; Shi et al., 2024; Ou et al., 2025; Zheng et al., 2025), decoding and sampling efficiency (Zheng et al., 2025; Kim et al., 2025; Zhao et al., 2025; Park et al., 2024; Liu et al., 2025), and large-scale model development (Nie et al., 2025; Zhu et al., 2025a; Song et al., 2025; Ye et al., 2025; Xie et al., 2025; Khanna et al., 2025). Together, these studies have substantially improved our understanding of dLLMs and have made diffusion-based language generation increasingly effective and scalable.
Our work is most closely related to this recent line of dLLMs, but differs in focus. Rather than designing a new corruption process or decoding heuristic, we develop a distillation framework for few-step generation based on distribution matching. Inspired by the continuous-time IKL perspective, our method is tailored to models so as to maintain tractable training and inference while improving generation efficiency and quality.
A.3 Related Work on Few-Step Diffusion Large Language Models
@@
To distill dLLMs into few-step generators for fast language generation, a natural approach is to extend the well-developed distillation techniques already established for continuous diffusion models to the discrete domain. However, in MDMs, all tokens are eventually mapped to the masked state, meaning the prior distribution collapses to the fully masked sequence. As a result, generation under MDMs is inherently stochastic, in contrast to the deterministic trajectories available in the continuous case. Consequently, the stochasticity in Masked Diffusion Models (MDMs) arises from the choice of the masking schedule; once a schedule is fixed, the reverse path is deterministic. This stands in sharp contrast to the inherently stochastic reverse dynamics of continuous diffusion. This determinism is problematic for distillation, as training a student model solely on these fixed trajectories risks mode collapse and a failure to capture the teacher’s generalization capabilities. This fundamental difference makes the extensive body of work on continuous-state diffusion distillation not directly applicable. To make this distinction concrete, we first show how the following methods address these challenges in discrete diffusion distillation, followed by a comparison of these representative approaches.
SDTT (Deschenaux and Gulcehre, 2025).
@@
SDTT progressively distills a multistep teacher diffusion into a fewer-step student model. Due to the absence of an ODE trajectory as in continuous cases, it matches the teacher-student distribution via KL divergence minimization. Similar to DUO (Sahoo et al., 2025), it needs to decrease the number of student steps for training stability gradually. Such a curriculum-based strategy can be inefficient and computationally costly.
DUO (Sahoo et al., 2025).
@@
Continuous-space consistency distillation (CD) requires probability flow ODE (PF-ODE) trajectories, while MDMs are inherently stochastic, making direct CD inapplicable. DUO overcomes this by connecting Gaussian diffusion with Uniform-state Diffusion Model (USDM), enabling CD in the discrete setting. By applying an $\operatorname*{arg\,max}$ to the latent vectors of Gaussian diffusion, continuous vectors are mapped into discrete one-hot tokens. It is proven that the resulting marginal distribution exactly matches that of a USDM under a suitably transformed noise schedule. Consequently, the evolution of a USDM can be described by an ODE, which makes CD feasible. However, performing CD on USDM requires multiple rounds of distillation and a carefully annealed discretization schedule for training stability. Specifically, the CD procedure starts from the USDM diffusion and gradually learns to map two points, $t$ and $t-\Delta t$, along the same ODE trajectory back to the same origin. Initially, since the diffusion model cannot take large steps without losing accuracy, $\Delta t$ must be kept small. Once the model learns to consistently align $t$ and $t-\Delta t$, the step size can be increased, eventually reaching the largest possible step $\Delta t=t$, i.e., the distilled model can jump large steps for few-step inference. A small step is safe but biased, while a large step is unbiased but unstable. This tradeoff is similar to the observation of CD in the continuous domain.
DiMO (Zhu et al., 2025b).
@@
DiMO distills multi-step MDMs into a one-step generator. The key idea is to augment the prior distribution: instead of restricting the initial state to a fully-masked sequence, DiMO samples a subset of tokens from the entire vocabulary. This relaxation enriches the prior, allowing the model to leverage partial information during generation and improving both efficiency and diversity of outputs. During training, DiMO employs an on-policy distillation strategy. The one-step generator is supervised to match the teacher’s conditional prediction distribution, not its final output distribution. Specifically, it uses the generator’s own one-step output to create a pseudo-intermediate state, and then minimizes the divergence between the student’s and the teacher’s predicted token distributions conditioned on this state. To avoid mode collapse and reduce mismatch with the teacher’s training distribution, DiMO introduces a token initialization strategy that mixes mask tokens with random tokens and adds Gaussian perturbations to embeddings.
@@
Although DiMO shows promise in text-to-image generation, it should be noted that extending it to our setting reveals a significant modality gap. DiMO is specialized for image codebook tokens characterized by 2D spatial dependencies and visual coherence. This differs fundamentally from pure language generation, which necessitates modeling 1D sequential dependencies and strict syntactic constraints. In our preliminary experiments, a direct adaptation of DiMO yielded comparable generative perplexity but significantly degraded diversity (entropy drops to 4.0), which suggests training instability and potential mode collapse under naive adaptation. Consequently, transferring architectural choices tuned for visual latents to the text domain is non-trivial, and we consider a principled, text-specific adaptation of DiMO to be a valuable area for future work.
Our advantages over three related methods.
Our work presents a straightforward and intuitive framework for distilling dLLMs by directly matching teacher and student distributions, addressing several key limitations of recent approaches.
Unlike trajectory-based methods (Sahoo et al., 2025; Deschenaux and Gulcehre, 2025) that rely on supervised trajectory matching (forcing the student to mimic the teacher’s logits along rigid, predetermined paths), DiDi-Instruct is a direct, single-stage process derived from Integral KL minimization. By reformulating distillation as a policy gradient problem, DiDi-Instruct avoids the heuristic mappings and expensive multi-stage training typical of trajectory-based methods. Crucially, this objective exhibits a natural explore-exploit mechanism: the stochastic sampling of intermediate states $t_{i}$ and $\mathrm{z}_{i}$ in (9) facilitates exploration of the discrete tokens, while the student’s few-step backward update promotes exploitation. This natural mechanism effectively mitigates the mode-collapse issues inherent in trajectory-matching baselines.
Furthermore, by aligning the student with the teacher’s entire stochastic process, our framework offers broad applicability to general dLLMs, in contrast to methods tailored for specific architectures, e.g., DUO (Sahoo et al., 2025) relies on USDM for duality. Finally, we provide a mathematically rigorous solution to the problem of non-differentiability in discrete spaces. This avoids the need for biased proxy-gradient estimators, whose performance in purely textual domains remains underexplored (Zhu et al., 2025b), and instead offers a principled path for distillation.
Appendix B Student Objective Derivation
Our goal is to train a masked-diffusion student whose forward-time marginals match those of a teacher across the entire time horizon. The training objective is the IKL divergence between the student and teacher forward marginals, denoted by $\mathbf{q}_{\nu}$ and $\mathbf{q}_{\theta}$. In the masked setting, the forward-time corruption kernel is a forward absorbing process, e.g.,